A father and son are building a round fence around their house. They have 16, 16ft. panels and a 10 ft. gate and a 4ft. gate (270 ft. total). They want to use a radius and mark the places to dig holes for each post that will support the panels, but they don't know how long the radius should be. Find the radius
Answer by Cromlix(4381)

If the total length of circular fence = 270 ft
this equals the circumference of a circle.
Circumference = Pi * diameter
Pi * diameter = 270 ft
Diameter = 270/Pi
Diameter = 85.944 (3 significant figures)
Diameter = 2 * radius
Radius = 85.944/2
Radius = 42.97 ft or rounded up 43 feet.
Hope this helps.
:-)